# spacing=nonfrench

### v1

@@ Abstract
Anti-malware systems rely on sandboxes, hooks, and telemetry pipelines, including collection agents, serializers, and database backends, to monitor program and system behavior. We show that these data-handling components constitute an exploitable attack surface that can lead to denial-of-analysis (DoA) states without disabling sensors or requiring elevated privileges. As a result, we present Telemetry Complexity Attacks (TCAs), a new class of vulnerabilities that exploit fundamental mismatches between unbounded collection mechanisms and bounded processing capabilities. Our method recursively spawns child processes to generate specially crafted, deeply nested, and oversized telemetry that stresses serialization and storage boundaries, as well as visualization layers, for example, JSON/BSON depth and size limits. Depending on the product, this leads to truncated or missing behavioral reports, rejected database inserts, serializer recursion and size errors, and unresponsive dashboards. In all of these cases, malicious activity is normally executed; however, depending on the examined solution, it is not recorded and/or not presented to the analysts. Therefore, instead of evading sensors, we break the pipeline that stores the data captured by the sensors.
@@ 2 Concept and methodology
Let us consider a trivial example of a malware that is executed in a malware sandbox. Its execution triggers some processes for which the malware sandbox would timestamp, keep track of the parent and children processes or even store their memory dumps. The natural way to represent this information is through a tree; however, if you need to share this information with others or store it in a database, this data structure might not be the most convenient way. This convenience may be easily achieved by using a JSON representation of the tree through nesting. Yet, arbitrary nesting is not natively supported by all libraries or databases.
@@ 2.1 Threat model and assumptions
We assume that an adversary can store an unprivileged user-mode binary on either (i) an endpoint that is instrumented by sandboxing or EDR/DFIR tooling, or (ii) at a cloud solution, by uploading the binary through common user interfaces. The binary recursively spawns child processes up to a chosen depth and rate. Execution occurs under typical analysis conditions (e.g., detonation VMs, lab hosts, or enterprise-managed endpoints) where process creation is permitted and environment variables propagate to child processes.
@@
This work does not aim to achieve persistence, privilege escalation, data exfiltration, or evasion of dynamic behavior capture. We do not claim cryptographic bypasses, supply-chain compromise, or exploitation of memory corruption. Likewise, we did not introduce obfuscation mechanisms that would completely blind the payload, but only the necessary ones to prevent EDRs from blocking execution from static analysis. As such, we only focus on introducing operational failures in telemetry pipelines using valid but manipulated inputs generated from the data sources they are programmed to collect.
@@ 3.1 Setup
The experiments followed a progressive strategy. We started with open-source sandboxing platforms to iterate quickly and observe internal failures. After validating the concept on those systems, we moved to open-source EDR and DFIR tools and locally deployed trial instances of commercial products. Finally, sanitized samples were submitted to public, cloud-based analysis services to validate behavior with online providers. All local testing was performed in isolated, offline laboratory networks using virtual machines, to prevent telemetry leakage and to ensure safe handling of test artifacts. Each sandbox or EDR instance was installed with default or documented configurations; where supported, we varied timeouts, process depth limits, memory dump options, and report compression settings to exercise different code paths. Tests were repeated multiple times to ensure reproducibility; observed failures were cross-validated across independent environments when possible. Logging and monitoring of host resource usage (CPU, memory, disk I/O) were performed for each run to correlate system load with observed telemetry failures. The isolation of the experimental environment from the internet was introduced to avoid the communication of telemetry to software vendors that would allow them to prevent the execution of our binaries by adding some static checks or patching their solution on the fly.
@@ 3.4 Payload obfuscation
To reduce the likelihood of early static detection from EDRs, the sample employs a multi-layer ‘‘onion’’ encryption scheme. Each runtime stage decrypts only the subsequent stage. This prevents the full payload or final command-and-control instruction from being visible during static inspection. The obfuscation is intentionally lightweight, sufficient to bypass basic static heuristics while keeping behavior reproducible across testbeds.
@@ 4.3 Undetected malicious behavior from malware sandboxes
In the case of some malware sandboxes, e.g., Bitdefender GravityZone Sandbox, Threat Zone, while there is a verdict, the solution does not manage to detect what the malware was doing and report it to the analyst. In fact, in some cases, the solution might not have broken, e.g., Threat Zone, but the verdict was that the sample was benign.
@@ 4.5 CAPEv2
Our tests resulted in a Denial-of-analysis in reporting/mongodb.py and reporting/jsondump.py in CAPEv2 that allows attackers who can submit samples to cause incomplete or missing behavioral analysis reports by generating deeply nested or oversized behavior data that trigger MongoDB BSON limits or orjson recursion errors when the sample executes in the sandbox. Note that many systems depend on CAPEv2 dynamic analysis, e.g., VirusTotal uses it as one of its sandboxes. This issue has been assigned CVE-2025-61301 https://nvd.nist.gov/vuln/detail/CVE-2025-61301.
@@ 4.6 Triage
Reproducible public detonations are available on Triage [^4] https://tria.ge/250822-qfystshk51, which we used to validate the behavior across the provider’s Windows images. Based on the Triage’s API, we know that the behaviour analysis report is a JSON, which, for the processes, uses a nested format. Therefore, we assume that the same process is followed when collecting the events, leading to a crash that truncates the output. We do not consider that the issue is in the presentation layer, as a score is fetched (1/10), and part of the process tree is reported.
@@
We cannot disclose the vendor name, product build identifiers, or specific commit references at this time, because the vendor requested an additional 30 day remediation window to deploy multiple patches across their systems. The vendor acknowledged the issue and is actively working on patching; we have submitted the report to MITRE for CVE assignment and are awaiting a CVE identifier. Per our coordinated disclosure agreement, details will remain private until the remediation window elapses. Therefore, for the remainder of the subsection we refer to the affected platform simply as <REDACTED>.
@@
We reproduced the issue on both public instances and locally cloned deployments. In the affected analyses, we regularly saw <REDACTED> and <REDACTED> grow to 45-50 MB, so the underlying collectors were clearly logging a lot of activity; however, the final behavioral report was empty. Attempting to open those reports often caused the web UI to stall or the browser to freeze, and in offline clones, the large report files and UI instability prevented access to otherwise completed analyses. Representative runs and measures artifacts are summarized in Table 3.
@@ 5.1 Root cause analysis and mitigation measures
Based on our findings, several practical steps can reduce denial-of-analysis risk across monitoring platforms and the impact of TCAs.
@@
The reporting mechanisms should also become more robust. For example, incremental rendering of dashboards and reports can prevent interface issues with oversized datasets and selectively provide truncated previews when needed. In most cases, not all information has to be displayed to the user at once; therefore, some information can be prioritized, hinting that further information exists, improving usability and stability. An assessment of how much information exists and how this can be properly loaded, given the memory limits, internal size, and recursion thresholds, can also facilitate this process. Thus, dashboards must render lazily, using, e.g., virtualized trees and pagination, and avoid attempts to materialize large hierarchies in the browser or server render threads. This can be facilitated by using streaming JSON/BSON parsers to avoid loading entire documents into memory, enabling partial consumption and checkpointing of long streams.
@@ 6 Related work
Dynamic analysis of binaries with and without the use of automated sandboxing is central to malware analysis. Prior work has shown various ways in which adversaries try to defeat these systems. Evasion of malware sandboxes usually aims at the identification of the sandbox environment through various means, such as timing attacks against virtualization, registry, network, or CPU artifact attacks, and human interaction detection [4]. For example, SandPrint showed that sandbox implementations may leak stable fingerprints that malware can detect and react to [25]. Timing and resource channels remain practical. Embedding proof-of-work into malware exposes timing asymmetries between bare-metal and scaled cloud sandboxes, reducing the effectiveness of online analysis services [17].
@@ 6.1 Attacks on EDRs
One of the main directions in this research line targets the hooking mechanisms used by EDRs to intercept user-mode or kernel-mode APIs. Several works exploit unhooking or direct system-call invocation to execute unmonitored code paths, typically by calling native Windows functions from ntdll.dll without passing through instrumented stubs [2, 9, 14]. Because many EDRs rely on user-land hooks, both attackers and defenders operate at the same privilege level, creating inherent race conditions, which at the end favor the attackers. Recent work, such as Cymulate’s Blindside technique, employed hardware breakpoints to selectively allow the loading of ntdll.dll while blocking other libraries. [^5] https://cymulate.com/blog/blindside-a-new-technique-for-edr-evasion-with-hardware-breakpoints/
@@ 6.2 Attacks on malware sandboxes
Contrary to EDRs, malware sandboxes allow malware to execute, as their goal is to actually observe what the malware does when it is executed. Therefore, malware tries to hide as much information as possible. Therefore, malware needs to identify the environment in which it operates and determine whether it will continue its operation. For instance, since the malware sandbox is a virtualized environment, malware tries to determine whether they are being executed in a virtual machine. Side information, e.g., names of devices or running processes, hardware identifiers, screen resolution, available memory, or even available threads and processors can easily give away the execution in a virtualized environment [5, 21, 12]. Similarly, a malware may try to find drivers, registry keys, or even folder names associated with virtual machines. Going a step further, malware may try to find elements of human interaction to determine whether it is executed in a malware sandbox. For instance, the uptime, the lack of keyboard and mouse interaction, the absence of browser history, or recent file activity can be used to infer the environment. Therefore, many malware would perform these checks before execution of the actual payload. Moreover, this creates a very fragile and sensitive environment where different configurations can provide radically different reports.
Given the resource and time constraints, some malware may opt to delay their payload execution [11]. This forced malware sandboxes to manipulate time to trigger their execution faster. As a result of this arms race, malware tries to measure the timing of specific functions to determine whether the elapsed time is aligned with the expected processing time. Evidently, time plays an important role in the dynamic analysis of malware [13].
@@ 6.3 Denial-of-Service attacks on security solutions
Petsios et al. presented SlowFuzz, a method to automatically synthesize inputs that trigger worst-case for several well-known algorithms, causing severe behavioral issues in widely deployed libraries [18]. For what is relevant in this work, many various zip parsers used in antivirus software whose decompression was significantly delayed. Other related work detected and repaired DoS-prone regex patterns, often used by cybersecurity solutions to detect strings or bytes associated with malicious activity (e.g., ReDoSHunter, RegexScalpel) and exposed vulnerabilities in non-backtracking matchers [15, 24, 16]. In content-inspection pipelines, adversarial inputs that induce super-linear or exponential matching can exhaust CPU budgets, delay processing, and cause packet or alert drops.

### v2

@@ Abstract
Anti-malware systems rely on sandboxes, hooks, and telemetry pipelines, including collection agents, serializers, and database backends, to monitor program and system behavior. We show that these data-handling components constitute an exploitable attack surface that can lead to denial-of-analysis (DoA) states without disabling sensors or requiring elevated privileges. As a result, we present Telemetry Complexity Attacks (TCAs), a new class of vulnerabilities that exploit fundamental mismatches between unbounded collection mechanisms and bounded processing capabilities. Our method recursively spawns child processes to generate specially crafted, deeply nested, and oversized telemetry that stresses serialization and storage boundaries, as well as visualization layers, for example, JSON/BSON depth and size limits. Depending on the product, this leads to various inconsistent results, such as truncated or missing behavioral reports, rejected database inserts, serializer recursion and size errors, and unresponsive dashboards. In the latter cases, depending on the solution, the malware under test is either not recorded and/or not presented to the analysts. Therefore, instead of evading sensors, we break the pipeline that stores the data captured by the sensors.
@@ 2 Concept and methodology
Let us consider a trivial example of a malware that is executed in a malware sandbox. Its execution triggers some processes for which the malware sandbox would timestamp, keep track of the parent and children processes, or even store their memory dumps. The natural way to represent this information is through a tree; however, if you need to share this information with others or store it in a database, this data structure might not be the most convenient way. This convenience may be easily achieved by using a JSON representation of the tree through nesting. Yet, arbitrary nesting is not natively supported by all libraries or databases.
@@ 2.1 Threat model and assumptions
We assume that an adversary can store an unprivileged user-mode binary on either (i) an endpoint that is instrumented by sandboxing or EDR/DFIR tooling, or (ii) at a cloud solution, by uploading the binary through common user interfaces. The binary recursively spawns child processes up to a chosen depth and rate. Execution occurs under typical analysis conditions (e.g., detonation VMs, lab hosts, or enterprise-managed endpoints), where process creation is permitted and environment variables are propagated to child processes.
@@
This work does not aim to achieve persistence, privilege escalation, data exfiltration, or evasion of dynamic behavior capture. We do not claim cryptographic bypasses, supply-chain compromise, or exploitation of memory corruption. Likewise, we did not introduce obfuscation mechanisms that would completely blind the payload, but only the necessary ones to prevent EDRs from blocking execution due to static analysis. As such, we only focus on introducing operational failures in telemetry pipelines using valid but manipulated inputs generated from the data sources they are designed to collect.
@@ 3.1 Setup
The experiments followed a progressive strategy. We started with open-source sandboxing platforms to iterate quickly and observe internal failures. After validating the concept on those systems, we moved to open-source EDR and DFIR tools and locally deployed trial instances of commercial products. Finally, sanitized samples were submitted to public, cloud-based analysis services to validate behavior with online providers. All local testing was performed in isolated, offline laboratory networks using virtual machines, to prevent telemetry leakage and to ensure safe handling of test artifacts. Each sandbox or EDR instance was installed with default or documented configurations; where supported, we varied timeouts, process depth limits, memory dump options, and report compression settings to exercise different code paths. Tests were repeated multiple times to ensure reproducibility; observed failures were cross-validated across independent environments when possible. Logging and monitoring of host resource usage (CPU, memory, disk I/O) were performed for each run to correlate system load with observed telemetry failures. The isolation of the experimental environment from the internet was introduced to prevent the communication of telemetry to software vendors, which would allow them to prevent the execution of our binaries by adding static checks or patching their solution on the fly.
@@ 3.4 Payload obfuscation
To reduce the likelihood of early static detection by EDRs, the sample employs a multi-layer ‘‘onion’’ encryption scheme. Each runtime stage decrypts only the subsequent stage. This prevents the full payload or final command-and-control instruction from being visible during static inspection. The obfuscation is intentionally lightweight, sufficient to bypass basic static heuristics while maintaining reproducible behavior across testbeds.
@@ 4.3 Undetected malicious behavior from malware sandboxes
In the case of some malware sandboxes, e.g., Bitdefender GravityZone Sandbox and Threat Zone, even though a verdict is reached, the solution fails to detect what the malware was doing and report it to the analyst. In fact, in some cases, the solution might not have broken, e.g., Threat Zone, but the verdict was that the sample was benign.
@@ 4.5 CAPEv2
Our tests resulted in a Denial-of-analysis in reporting/mongodb.py and reporting/jsondump.py in CAPEv2 that allows attackers who can submit samples to cause incomplete or missing behavioral analysis reports by generating deeply nested or oversized behavior data that trigger MongoDB BSON limits or orjson recursion errors when the sample executes in the sandbox. Note that many systems rely on CAPEv2 dynamic analysis, e.g., VirusTotal uses it as one of its sandboxes. This issue has been assigned CVE-2025-61301 https://nvd.nist.gov/vuln/detail/CVE-2025-61301.
@@ 4.6 Triage
Reproducible public detonations are available on Triage [^4] https://tria.ge/250822-qfystshk51, which we used to validate the behavior across the provider’s Windows images. Based on the Triage’s API, we know that the behaviour analysis report is a JSON, which, for the processes, uses a nested format. Therefore, we assume that the same process is followed when collecting the events, leading to a crash that truncates the output. We do not consider the issue to be in the presentation layer, as a score is fetched (1/10) and part of the process tree is reported.
@@
We cannot disclose the vendor name, product build identifiers, or specific commit references at this time because the vendor requested an additional 30-day remediation window to deploy multiple patches across their systems. The vendor acknowledged the issue and is actively working on patching; we have submitted the report to MITRE for CVE assignment and are awaiting a CVE identifier. As per our coordinated disclosure agreement, details will remain private until the remediation window has elapsed. Therefore, for the remainder of the subsection, we refer to the affected platform simply as <REDACTED>.
@@
We reproduced the issue on both public instances and locally cloned deployments. In the affected analyses, we regularly saw <REDACTED> and <REDACTED> grew to 45-50 MB, so the underlying collectors were clearly logging a lot of activity; however, the final behavioral report was empty. Attempting to open those reports often caused the web UI to stall or the browser to freeze, and in offline clones, the large report files and UI instability prevented access to otherwise completed analyses. Representative runs and measures artifacts are summarized in Table 3.
@@ 5.1 Root cause analysis and mitigation measures
Based on our findings, several practical steps can reduce the risk of denial-of-analysis across monitoring platforms and the impact of TCAs.
@@
The reporting mechanisms should also become more robust. For example, incremental rendering of dashboards and reports can prevent interface issues with oversized datasets and selectively provide truncated previews when needed. In most cases, not all information needs to be displayed to the user at once; therefore, some information can be prioritized, hinting that further information exists, improving usability and stability. An assessment of how much information exists and how this can be properly loaded, given the memory limits, internal size, and recursion thresholds, can also facilitate this process. Thus, dashboards must render lazily, using, e.g., virtualized trees and pagination, and avoid attempts to materialize large hierarchies in the browser or server render threads. This can be facilitated by using streaming JSON/BSON parsers to avoid loading entire documents into memory, enabling partial consumption and checkpointing of long streams.
@@ 6 Related work
Dynamic analysis of binaries with and without the use of automated sandboxing is central to malware analysis. Prior work has demonstrated various methods by which adversaries attempt to compromise these systems. Evasion of malware sandboxes typically aims to identify the sandbox environment through multiple means, such as timing attacks against virtualization, registry, network, or CPU artifact attacks, as well as detection of human interaction [4]. For example, SandPrint showed that sandbox implementations may leak stable fingerprints that malware can detect and react to [25]. Timing and resource channels remain practical. Embedding proof-of-work into malware exposes timing asymmetries between bare-metal and scaled cloud sandboxes, reducing the effectiveness of online analysis services [17].
@@ 6.1 Attacks on EDRs
One of the main directions in this research line targets the hooking mechanisms used by EDRs to intercept user-mode or kernel-mode APIs. Several works exploit unhooking or direct system-call invocation to execute unmonitored code paths, typically by calling native Windows functions from ntdll.dll without passing through instrumented stubs [2, 9, 14]. Because many EDRs rely on user-land hooks, both attackers and defenders operate at the same privilege level, creating inherent race conditions, which, in the end, favor the attackers. Recent work, such as Cymulate’s Blindside technique, employed hardware breakpoints to selectively allow the loading of ntdll.dll while blocking other libraries. [^5] https://cymulate.com/blog/blindside-a-new-technique-for-edr-evasion-with-hardware-breakpoints/
@@ 6.2 Attacks on malware sandboxes
Contrary to EDRs, malware sandboxes allow malware to execute, as their goal is to actually observe what the malware does when it is executed. Therefore, malware tries to hide as much information as possible. Therefore, malware needs to identify the environment in which it operates and determine whether it will continue its operation. For instance, since the malware sandbox is a virtualized environment, malware attempts to determine whether it is being executed in a virtual machine. Side information, e.g., names of devices or running processes, hardware identifiers, screen resolution, available memory, or even available threads and processors can easily give away the execution in a virtualized environment [5, 21, 12]. Similarly, malware may try to find drivers, registry keys, or even folder names associated with virtual machines. Going a step further, malware may try to find elements of human interaction to determine whether it is executed in a malware sandbox. For instance, factors such as uptime, the lack of keyboard and mouse interaction, the absence of browser history, or recent file activity can be used to infer the environment. Therefore, many malware would perform these checks before executing the actual payload. Moreover, this creates a very fragile and sensitive environment where different configurations can provide radically different reports.
Given resource and time constraints, some malware may opt to delay the execution of their payload [11]. This forced malware sandboxes to manipulate time to trigger their execution faster. As a result of this arms race, malware tries to measure the timing of specific functions to determine whether the elapsed time is aligned with the expected processing time. Evidently, time plays a vital role in the dynamic analysis of malware [13].
@@ 6.3 Denial-of-Service attacks on security solutions
Petsios et al. presented SlowFuzz, a method to automatically synthesize inputs that trigger worst-case scenarios for several well-known algorithms, causing severe behavioral issues in widely deployed libraries [18]. For what is relevant in this work, many various zip parsers are used in antivirus software whose decompression was significantly delayed. Other related work detected and repaired DoS-prone regex patterns, often used by cybersecurity solutions to detect strings or bytes associated with malicious activity (e.g., ReDoSHunter, RegexScalpel) and exposed vulnerabilities in non-backtracking matchers [15, 24, 16]. In content-inspection pipelines, adversarial inputs that induce super-linear or exponential matching can exhaust CPU budgets, delay processing, and cause packet or alert drops.